OpenStudy (anonymous):

Solve algebraically (x-5)^2(x+2)<0

OpenStudy (anonymous):

\[\left( x-5 \right)^2\left( x+2 \right)<0\] \[\left( x-5 \right)^2 always~positive.\] only possibility is x+2 <0 x<-2